A box contains 28 transitors, 5 of which are defective. If 5 are sold at random, find the probability that at least 1 is defective. Express your answer as a fraction.
Answer by stanbon(75887)

A box contains 28 transitors, 5 of which are defective. If 5 are sold at random, find the probability that at least 1 is defective. Express your answer as a fraction.
--------
P(>=1 defective) = 1 - P(none defective)
---
= 1 - 28C5/33C5
---
= 1 - 0.4141
---
= 0.5859
=================
Cheers,
Stan H.